OpenStudy (anonymous):

y=log[(10-x)/(x+2)] , find y' .

OpenStudy (anonymous):

\[y=log[\frac{10-x}{x+2}]\] \[y=log[10-x]-log[x+2]\] \[y\prime=\frac{-1}{10-x}-\frac{1}{x+2}\]